OpenStudy (mathmath333):

If a card is picked at random from a pack of 52 cards. Find the probablity that it is 'a spade ' or 'a king' or 'a queen'?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ If a card is picked at random from a pack}\hspace{.33em}\\~\\ & \normalsize \text{ of 52 cards. Find the probablity that it is }\hspace{.33em}\\~\\ & \normalsize \text{ 'a spade ' or 'a king' or 'a queen'}\hspace{.33em}\\~\\ & a.)\ \dfrac{21}{52} \hspace{.33em}\\~\\ & b.)\ \dfrac{5}{13} \hspace{.33em}\\~\\ & c.)\ \dfrac{19}{52} \hspace{.33em}\\~\\ & d.)\ \dfrac{15}{52} \hspace{.33em}\\~\\ \end{align}}\)

Nnesha (nnesha):

add number of queen, king and spades cards divide by total number of cards

Nnesha (nnesha):

i think XD

OpenStudy (jango_in_dtown):

P(A+B+C)=P(A+(B+C))=P(A)+P(B+C)-P(A)P(B+C)=P(A)+P(B)+P(C)-P(BC)-P(AB)-P(CA)+P(ABC).. So P(A+B+C)=P(A)+P(B)+P(C)-P(AB)-P(BC)-P(CA)+P(ABC)

OpenStudy (jango_in_dtown):

You need to use the above formula to solve this problem

Nnesha (nnesha):

i think there are 13 spades 4 king and 4 queen 4+4+13 /52 lol is this correct ?

Nnesha (nnesha):

oh okay nvm then..

OpenStudy (jango_in_dtown):

Let A denote the event that a spade is picked, B denote the event that a king is picked and C denote the event that a queen is picked,,, then our required probability is P(A+B+C)

OpenStudy (mathstudent55):

You must subtract cards in common.

OpenStudy (jango_in_dtown):

P(A)=13/52=1/4 P(B)=4/52=1/13 P(C)=4/52=1/13 P(AB)=1/52 P(BC)=0 P(CA)=1/52 P(ABC)=0

Directrix (directrix):

13 spades, 3 kings which are not spades, 3 queens that are not spades P(spade or king or queen) = 13/52 + 3/52 + 3/52 =

OpenStudy (mathstudent55):

spade: 13 cards (including a king of spades and a queen of spades) king: 3 cards (other than king of spades already counted above) queen: 3 cards (other than queen of spades already counted above) Total number of cards of interest: 13 + 3 + 3 = 19 Total number of cards in deck: 52 P(spade or king or queen) = 19/52

OpenStudy (jango_in_dtown):

Now just plug in and you will get required probability=(1/4)+(1/13)+(1/13)-(1/52)-(1/52) =(13+4+4-1-1)/52=19/52

OpenStudy (mathmath333):

i like that formula i forgot of p(a+b+c)

OpenStudy (jango_in_dtown):

I derived the fomula , you may check

Nnesha (nnesha):

ah 3 king and 3 queens lack of knowledge abt cards \,,/(>.<)\,,/ :D

OpenStudy (trojanpoem):

I used google and it thinks that there is 4 queens , 4 kings.

Directrix (directrix):

@Nnesha 4 kings and 4 queens in the deck 3 kings and 3 queens in the deck are not spades. See graphic below.

OpenStudy (trojanpoem):

a spade ' or 'a king' or 'a queen ? 13, 4 , 4 ?

Nnesha (nnesha):

aw i see! thanks i'll keep that n my mind..\o^_^o/